# Breaking Neural Network Scaling Laws with Modularity

### v1

@@ 1 Introduction
Modular neural network (NN) architectures have achieved impressive results in a variety of domains ranging from visual question answering (VQA) (Andreas et al., 2016a; b; Hu et al., 2017; Johnson et al., 2017; Yi et al., 2018; Kim et al., 2019), reinforcement learning (Goyal et al., 2021; Madan et al., 2021), robotics (Alet et al., 2018b; Pathak et al., 2019; Yang et al., 2020) and natural language processing for which modular architectures based on attention (Bahdanau et al., 2015) are standard. Modular NNs are thought to be better generalized by facilitating combinatorial generalization, a phenomenon where a learning system recombines previously learned components in novel ways to generalize to unseen task inputs (Alet et al., 2018b; D’Amario et al., 2021; Mittal et al., 2022a; b; Jarvis et al., 2023). Yet, a fundamental understanding of why modularity benefits generalization is lacking.
@@ 3.1 Model Setup
We defer the full details of our theoretical model setup to App A. Here we present an overview: we consider a regression task with input $x\in\mathbb{R}^{m}$ and a feature matrix $\varphi(x)\in\mathbb{R}^{d\times P}$, where $m$ is input dimensionality, $d$ is output dimensionality and $P$ is the number of features. We assume that over the data distribution, the features are distributed i.i.d. from a unit Gaussian: $\varphi(x)_{i,j}\sim\mathcal{N}(0,1)$. We consider the limit when $P\to\infty$. Suppose that our regression target function $y:\mathbb{R}^{m}\to\mathbb{R}^{d}$ is constructed linearly from $\varphi(x)$: $y(x)=\varphi(x)W$. Assume $\mathbb{E}[W]=0$ and $\mathbb{E}[WW^{T}]=\Lambda$, where $\Lambda$ is diagonal and $\operatorname{Tr}(\Lambda)$ is finite. Suppose we aim to approximate $y(x)$ using a model $\hat{y}(x)$ constructed as follows: $\hat{y}(x)=\varphi(x)\begin{bmatrix}I\\ 0\end{bmatrix}\theta$, where $\theta\in\mathbb{R}^{p\times 1}$ are model parameters and $p$ is the number of parameters. This corresponds to the model only being able to control $p$ of the $P$ true underlying parameters in the construction of $y$. We decompose $\Lambda$ blockwise as: $\Lambda=\begin{bmatrix}\Lambda_{1}&0\\ 0&\Lambda_{2}\end{bmatrix}$ where $\Lambda_{1}\in\mathbb{R}^{p\times p}$ and $\Lambda_{2}\in\mathbb{R}^{(P-p)\times(P-p)}$. To capture the dependence of the target function on the input dimensionality $m$, we parameterize $\Lambda$ as having individual elements $\lambda_{i}=c\left[i^{-\Omega^{-m}}-(i+1)^{-\Omega^{-m}}\right]$. This corresponds to the number of effective dimensions of variation of $W$ scaling exponentially with $m$, which is consistent with prior work (McRae et al., 2020). We also produce versions of our theoretical results without setting a specific form for $\lambda_{i}$. Finally, we consider learning $\theta$ as the minimum norm interpolating solution.
@@ 3.2 Theoretical Properties
Please see App B for a proof and App D for more details on $F(n,p)$. Under general $\lambda_{i}$, training and test error grow with $Tr(\Lambda_{2})$; the specific rate at which they grow or shrink with parameters $m$ and $p$ depends on how rapidly $Tr(\Lambda_{2})$ grows with $m$ and shrinks with $p$. Under the specific parameterization for $\lambda_{i}$ in Eqn 19, Fig 1 plots the value of the training and test set error for varying $n$ and number of parameters, holding $d=1$. Observe that there is a clear interpolation threshold at $p=n$ where the training loss becomes zero; for $p\geq n$ the model has sufficient capacity to perfectly interpolate the training set. The training loss is positive and increases with $n$ in the underparameterized regime ($p<n$) since the model lacks the capacity to fit increasing amounts of data $n$. At the $p=n$ threshold, the test loss dramatically increases, then decreases as $p$ increases beyond $n$. This is consistent with empirically observed behavior of overparameterized NNs (Belkin et al., 2019). Similarly, test loss decreases as $n$ increases beyond $p$.
@@ 4.1 Sample Complexity of Modular Learning
Recall in Sec 3, modeling a NN as a linear function of its parameters successfully captured its generalization properties. We aim to use this model to demonstrate the improved generalization of modular learning. For analytical traceability, we restrict our analysis to a specific modular setting that captures crucial aspects of many real-world modular learning scenarios. In practical settings, modules often handle low-dimensional inputs, such as attention maps in Andreas et al. (2016b). As such, we assume modules receive projected versions of the task input $x\in\mathbb{R}^{m}$. Our theoretical analysis will assume linear projections, but our method and experiments are also applied to non-linear projections. Moreover, we assume that the module outputs are summed to produce a final output, a feature of architectures such as Mixture of Experts (Shazeer et al., 2017). Specifically, consider a modular NN constructed as a linear combination of general NNs (each constituting a module) of low-dimensional projections of the input:
@@
where $\hat{U}_{j}\in\mathbb{R}^{m\times b}$ is a linear projection, $\hat{y}_{j}:\mathbb{R}^{b}\to\mathbb{R}^{d}$ is a NN. We will assume that $b$ is small, creating a bottleneck to each module’s input. We normalize by $\frac{1}{\sqrt{K}}$ to make the scale of $\hat{y}(x)$ invariant to $K$ (treating each term $\hat{y}_{j}(\hat{U}_{j}^{T}x)$ as independent, the sum of the terms has variance $O(K)$, so dividing my $\sqrt{K}$ makes their variance constant). Note that each module $\hat{y}_{j}$ is itself a monolithic NN with arbitary architecture; we do not restrict the form of the modules themselves. Using the linearity assumption of Eqn 18 (namely, the model is a linear function of parameters), we may model this as:
@@ 4.2 Modular Learning Rule
where $y_{j}$ are functions that depend on a potentially nonlinear projection of the input $x$ as represented by $y_{j}$ depending on both module projections $U_{j}$ and inputs $x$. Observe that this generalizes the linear projections considered before (in Eqn 6). Suppose we aim to model the target function by approximating the $U_{j}$ with $\hat{U}_{j}$ and the $y_{j}$ with $\hat{y}_{j}$ (parameterized as a neural network). We propose a kernel-based rule to learn the initializations of $\hat{U}_{j}$ from the training data; this allows us to efficiently learn the modules $\hat{y}_{j}$. Assume we are provided a set of training data $y(X)\in\mathbb{R}^{dn\times 1}$. Given the modular structure, we first aim to approximate the data as:
@@
where $\mathbf{K}\in\mathbb{R}^{dn\times dn}$ is a kernel matrix applied to the data corresponding to the following kernel: $\mathbf{K}(x_{1},x_{2})=\sum_{i=1}^{K}\varphi(x_{1}^{T};\hat{U}_{i})\varphi(x_{2}^{T};\hat{U}_{i})^{T}=\kappa(x_{1},x_{2};\hat{U}_{i})$, where $\kappa(x_{1},x_{2};\hat{U}_{i})$ is a module-conditional kernel between $x_{1}$ and $x_{2}$. Experimentally, we tailor $\kappa$ to the modular structure of the problem we consider. Note that the above analysis only applies when $\phi$ is sufficiently expressive (i.e. $pK>dn$), which is a natural assumption for typically overparameterized models like neural networks. When models do not satisfy this assumption, Eqn 15 yields a solution minimizing the (generally nonzero) error between the predicted and true $y(X)$. Importantly, in this case, minimizing the norm of $\theta$ with respect to $\hat{U}_{i}$ may not necessarily yield a lower prediction error.
App E describes the specific choice of $\kappa$. Alg 1 shows the full procedure to find a single module projection $\hat{U}_{i}$; each step of the algorithm simply applies gradient on Eqn 16 with respect to $\hat{U}_{i}$. We repeat this procedure $K$ times with different random initializations to find the initial values of all $K$ module projections in our architecture. Then, we train all module parameters (including the $\hat{U}_{i}$) via gradient descent on the task loss. We stress that our approach is applicable to a fairly general set of modular architectures of the form $\sum_{j}\hat{y}_{j}(x;\hat{U}_{j})$: it does not restrict modules to receive only linear projections of inputs, and, moreover, does not restrict the form of the modules.
@@ 4.3 Experimental Results
So far, we have considered modules ($y_{j}$ in Eqn 13) for which the input is a linear function of both $U_{j}$ and $x$. Next, we consider a nonlinear variant of the sine wave regression task in which modules are a function of $||u_{j}-x||_{2}$, where $u_{j}$ are the vector module projection directions, which is nonlinear in both $u_{j}$ and $x$. The modular architecture is constructed non-linearly as: $\frac{1}{\sqrt{K}}\sum_{j=1}^{K}\hat{y}_{j}(||\hat{u}_{j}-x||_{2})$; see App E for further details. Tab 2 illustrates that our method significantly outperforms the baselines, indicating that our method extends to nonlinear settings as well.
@@ 5 Discussion
In pursuit of explicit, non-asymptotic expressions for generalization error in modular and monolithic NNs, we make strong theoretical assumptions consistent with previous literature, which we hope can be further relaxed in future work. Moreover, our results apply to a specific form of modularity that captures structures in common real-world modular architectures but is not fully general. Notably, our theory considers linear module projections (although our method is also applied to nonlinear projections), and both the theory and the experiments assume that the model output is the sum of module outputs. We expect that future analyses can demonstrate the benefits of modularity more widely. For example, routing mechanisms are a popular type of modularity in which modules are flexibly composed or combined based on a routing network. We expect analyses similar to ours to show that, to the extent that routing mechanisms allow process low-dimensional inputs rather than the full task input, they also generalize better. Finally, we find that while the theory predicts a task-dimension-independent sample complexity for modular NNs, empirically we do not eliminate this dependence due to the difficulty of optimizing modular NNs in high dimensions. Nevertheless, our learning rule significantly eases this challenge.

### v2

@@ 1 Introduction
Modular neural network (NN) architectures have achieved impressive results in a variety of domains ranging from visual question answering (VQA) (Andreas et al., 2016a; b; Hu et al., 2017; Johnson et al., 2017; Yi et al., 2018; Kim et al., 2019), reinforcement learning (Goyal et al., 2021; Madan et al., 2021), robotics (Alet et al., 2018b; Pathak et al., 2019; Yang et al., 2020) and natural language processing for which modular architectures based on attention (Bahdanau et al., 2015) are standard. Modular NNs are thought to be better generalized by facilitating combinatorial generalization, a phenomenon where a learning system recombines previously learned components in novel ways to generalize to unseen inputs (Alet et al., 2018b; D’Amario et al., 2021; Mittal et al., 2022a; b; Jarvis et al., 2023). Yet, we lack a fundamental understanding of why modularity benefits generalization.
@@ 3.1 Model Setup
We defer the full details of our theoretical model setup to App A. Here we present an overview: we consider a regression task with input $x\in\mathbb{R}^{m}$ and a feature matrix $\varphi(x)\in\mathbb{R}^{d\times P}$ such that over the data distribution, the features are distributed i.i.d. from a unit Gaussian: $\varphi(x)_{i,j}\sim\mathcal{N}(0,1)$; no other assumptions are made about $\varphi$. We consider the limit when $P\to\infty$. Suppose that our regression target function $y:\mathbb{R}^{m}\to\mathbb{R}^{d}$ (mapping from $x$ to a $d$ dimensional output) is constructed linearly from $\varphi(x)$:
$y(x)=\varphi(x)W,$ (1)
where $W\in\mathbb{R}^{P\times 1}$. To accommodate multidimensional outputs, note that we shape input features $\varphi(x)$ as a matrix (with one row for each output dimension) and parameters $W$ as a vector. This choice is more general than parameterizing each output dimension independently (which can be captured as a special case of our approach) and , moreover, aligns with prior theoretical literature (Jacot et al., 2018). Assume $\mathbb{E}[W]=0$ and $\mathbb{E}[WW^{T}]=\Lambda$, where $\Lambda$ is diagonal and $\operatorname{Tr}(\Lambda)$ is finite. Suppose we aim to approximate $y(x)$ using a model $\hat{y}(x)$ constructed as follows: $\hat{y}(x)=\varphi(x)\begin{bmatrix}I\\ 0\end{bmatrix}\theta$, where $\theta\in\mathbb{R}^{p\times 1}$ are model parameters and $p$ is the number of parameters. This corresponds to the model only being able to control $p$ of the $P$ true underlying parameters in the construction of $y$. We decompose $\Lambda$ blockwise as: $\Lambda=\begin{bmatrix}\Lambda_{1}&0\\ 0&\Lambda_{2}\end{bmatrix}$ where $\Lambda_{1}\in\mathbb{R}^{p\times p}$ and $\Lambda_{2}\in\mathbb{R}^{(P-p)\times(P-p)}$. To capture the dependence of the target function on the input dimensionality $m$, we parameterize $\Lambda$ as having individual elements $\lambda_{i}=c\left[i^{-\Omega^{-m}}-(i+1)^{-\Omega^{-m}}\right]$. This corresponds to the number of effective dimensions of variation of $W$ scaling exponentially with $m$, which is consistent with prior work (McRae et al., 2020). We also produce versions of our theoretical results without setting a specific form for $\lambda_{i}$. Finally, we consider learning $\theta$ as the minimum norm interpolating solution.
@@ 3.2 Theoretical Properties
Please see App B for a proof and App D for more details on $F(n,p)$. Under general $\lambda_{i}$, training and test error grow with $Tr(\Lambda_{2})$; the specific rate at which they grow or shrink with parameters $m$ and $p$ depends on how rapidly $Tr(\Lambda_{2})$ grows with $m$ and shrinks with $p$. Under the specific parameterization for $\lambda_{i}$ described above, Fig 1 plots the value of the training and test set error for varying $n$ and number of parameters, holding $d=1$. Observe that there is a clear interpolation threshold at $p=n$ where the training loss becomes zero; for $p\geq n$ the model has sufficient capacity to perfectly interpolate the training set. The training loss is positive and increases with $n$ in the underparameterized regime ($p<n$) since the model lacks the capacity to fit increasing amounts of data $n$. At the $p=n$ threshold, the test loss dramatically increases, then decreases as $p$ increases beyond $n$. This is consistent with empirically observed behavior of overparameterized NNs (Belkin et al., 2019). Similarly, test loss decreases as $n$ increases beyond $p$.
@@ 4.1 Sample Complexity of Modular Learning
Recall in Sec 3, modeling a NN as a linear function of its parameters successfully captured its generalization properties. We aim to use this model to demonstrate the improved generalization of modular learning. For analytical traceability, we restrict our analysis to a specific modular setting that captures crucial aspects of many real-world modular learning scenarios. In practical settings, modules often handle low-dimensional inputs, such as attention maps in Andreas et al. (2016b). As such, we assume modules receive projected versions of the task input $x\in\mathbb{R}^{m}$. Our theoretical analysis will assume linear projections, but our method and experiments are also applied to non-linear projections. Moreover, we assume that the module outputs are summed to produce a final output, a feature of architectures such as Mixture of Experts (Jacobs et al., 1991; Shazeer et al., 2017). Specifically, consider a modular NN constructed as a linear combination of general NNs (each constituting a module) of low-dimensional projections of the input:
@@
where $\hat{U}_{j}\in\mathbb{R}^{m\times b}$ is a linear projection, $\hat{y}_{j}:\mathbb{R}^{b}\to\mathbb{R}^{d}$ is a NN. We will assume that $b$ is small, creating a bottleneck to each module’s input. We normalize by $\frac{1}{\sqrt{K}}$ to make the scale of $\hat{y}(x)$ invariant to $K$ (treating each term $\hat{y}_{j}(\hat{U}_{j}^{T}x)$ as independent, the sum of the terms has variance $O(K)$, so dividing my $\sqrt{K}$ makes their variance constant). Note that each module $\hat{y}_{j}$ is itself a monolithic NN with arbitrary architecture; we do not restrict the form of the modules themselves. Assuming that the model is a linear function of its parameters, we may model this as:
@@ 4.2 Modular Learning Rule
where $y_{j}$ are functions that depend on a potentially nonlinear projection of the input $x$ as represented by $y_{j}$ depending on both module projections $U_{j}$ and inputs $x$. Observe that this generalizes the linear projections considered before (in Eqn 7). Suppose we aim to model the target function by approximating the $U_{j}$ with $\hat{U}_{j}$ and the $y_{j}$ with $\hat{y}_{j}$ (parameterized as a neural network). We propose a kernel-based rule to learn the initializations of $\hat{U}_{j}$ from the training data; this allows us to efficiently learn the modules $\hat{y}_{j}$. Assume we are provided a set of training data $y(X)\in\mathbb{R}^{dn\times 1}$. Given the modular structure, we first aim to approximate the data as:
@@
where $\mathbf{K}\in\mathbb{R}^{dn\times dn}$ is a kernel matrix applied to the data corresponding to the following kernel: $\mathbf{K}(x_{1},x_{2})=\sum_{i=1}^{K}\varphi(x_{1}^{T};\hat{U}_{i})\varphi(x_{2}^{T};\hat{U}_{i})^{T}=\kappa(x_{1},x_{2};\hat{U}_{i})$, where $\kappa(x_{1},x_{2};\hat{U}_{i})$ is a module-conditional kernel between $x_{1}$ and $x_{2}$. Experimentally, we tailor $\kappa$ to the modular structure of the problem we consider. Note that the above analysis only applies when $\phi$ is sufficiently expressive (i.e. $pK>dn$), which is a natural assumption for typically overparameterized models like neural networks. When models do not satisfy this assumption, Eqn 16 yields a solution minimizing the (generally nonzero) error between the predicted and true $y(X)$. Importantly, in this case, minimizing the norm of $\theta$ with respect to $\hat{U}_{i}$ may not necessarily yield a lower prediction error.
App E describes the specific choice of $\kappa$. Alg 1 shows the full procedure to find a single module projection $\hat{U}_{i}$; each step of the algorithm simply applies gradient on Eqn 17 with respect to $\hat{U}_{i}$. We repeat this procedure $K$ times with different random initializations to find the initial values of all $K$ module projections in our architecture. Then, we train all module parameters (including the $\hat{U}_{i}$) via gradient descent on the task loss. We stress that our approach is applicable to a fairly general set of modular architectures of the form $\sum_{j}\hat{y}_{j}(x;\hat{U}_{j})$: it does not restrict modules to receive only linear projections of inputs, and, moreover, does not restrict the form of the modules.
@@ 4.3 Experimental Results
So far, we have considered modules ($y_{j}$ in Eqn 14) for which the input is a linear function of both $U_{j}$ and $x$. Next, we consider a nonlinear variant of the sine wave regression task in which modules are a function of $||u_{j}-x||_{2}$, where $u_{j}$ are the vector module projection directions, which is nonlinear in both $u_{j}$ and $x$. The modular architecture is constructed non-linearly as: $\frac{1}{\sqrt{K}}\sum_{j=1}^{K}\hat{y}_{j}(||\hat{u}_{j}-x||_{2})$; see App E for further details. Tab 2 illustrates that our method significantly outperforms the baselines, indicating that our method extends to nonlinear settings as well.
@@ 5 Discussion
In pursuit of explicit, non-asymptotic expressions for generalization error in modular and monolithic NNs, we make strong theoretical assumptions consistent with previous literature, which we hope can be further relaxed in future work. Moreover, our results apply to a specific form of modularity that captures structures in common real-world modular architectures but is not fully general. Notably, our theory considers linear module projections (although our method is also applied to nonlinear projections), and both the theory and the experiments assume that the model output is the sum of module outputs. We expect that future analyses can demonstrate the benefits of modularity more widely. For example, routing mechanisms are a popular type of modularity in which modules are flexibly composed or combined based on a routing network. We expect analyses similar to ours to show that, to the extent that routing mechanisms allow process low-dimensional inputs rather than the full task input, they also generalize better. Similar theory may also explain the generalization benefits of self-attention based architectures, which may learn modular substructures. Finally, we find that while the theory predicts a task-dimension-independent sample complexity for modular NNs, empirically we do not eliminate this dependence due to the difficulty of optimizing modular NNs in high dimensions. Nevertheless, our learning rule significantly eases this challenge. Further, our results suggest that more focus should be placed on optimization strategies for modern modular architectures beyond naively applying gradient descent: this could unlock further generalization benefits of commonly used modular motifs (such as routing mechanisms, attention, MoE etc.).